fill in the blank with one of the digits from 1 through 9 (using each digit only once) to make a true statement. using the following functions / then x then+ then x then x then / then + then x that = 100 (___ / ___ x ___ + ___ x ___ x ___ / ___ + ___ x ___ =100)


Answer by Edwin McCravy(20086)

Here are all 192 solutions:

  1.   1/2*6+5*7*8/4+3*9 = 100
  2.   1/2*6+5*7*8/4+9*3 = 100
  3.   1/2*6+5*8*7/4+3*9 = 100
  4.   1/2*6+5*8*7/4+9*3 = 100
  5.   1/2*6+7*5*8/4+3*9 = 100
  6.   1/2*6+7*5*8/4+9*3 = 100
  7.   1/2*6+7*8*5/4+3*9 = 100
  8.   1/2*6+7*8*5/4+9*3 = 100
  9.   1/2*6+8*5*7/4+3*9 = 100
 10.   1/2*6+8*5*7/4+9*3 = 100
 11.   1/2*6+8*7*5/4+3*9 = 100
 12.   1/2*6+8*7*5/4+9*3 = 100
 13.   4/1*8+2*3*5/6+7*9 = 100
 14.   4/1*8+2*3*5/6+9*7 = 100
 15.   4/1*8+2*5*3/6+7*9 = 100
 16.   4/1*8+2*5*3/6+9*7 = 100
 17.   4/1*8+3*2*5/6+7*9 = 100
 18.   4/1*8+3*2*5/6+9*7 = 100
 19.   4/1*8+3*5*2/6+7*9 = 100
 20.   4/1*8+3*5*2/6+9*7 = 100
 21.   4/1*8+5*2*3/6+7*9 = 100
 22.   4/1*8+5*2*3/6+9*7 = 100
 23.   4/1*8+5*3*2/6+7*9 = 100
 24.   4/1*8+5*3*2/6+9*7 = 100
 25.   4/6*9+2*5*7/1+3*8 = 100
 26.   4/6*9+2*5*7/1+8*3 = 100
 27.   4/6*9+2*7*5/1+3*8 = 100
 28.   4/6*9+2*7*5/1+8*3 = 100
 29.   4/6*9+5*2*7/1+3*8 = 100
 30.   4/6*9+5*2*7/1+8*3 = 100
 31.   4/6*9+5*7*2/1+3*8 = 100
 32.   4/6*9+5*7*2/1+8*3 = 100
 33.   4/6*9+7*2*5/1+3*8 = 100
 34.   4/6*9+7*2*5/1+8*3 = 100
 35.   4/6*9+7*5*2/1+3*8 = 100
 36.   4/6*9+7*5*2/1+8*3 = 100
 37.   4/8*6+2*5*7/1+3*9 = 100
 38.   4/8*6+2*5*7/1+9*3 = 100
 39.   4/8*6+2*7*5/1+3*9 = 100
 40.   4/8*6+2*7*5/1+9*3 = 100
 41.   4/8*6+5*2*7/1+3*9 = 100
 42.   4/8*6+5*2*7/1+9*3 = 100
 43.   4/8*6+5*7*2/1+3*9 = 100
 44.   4/8*6+5*7*2/1+9*3 = 100
 45.   4/8*6+7*2*5/1+3*9 = 100
 46.   4/8*6+7*2*5/1+9*3 = 100
 47.   4/8*6+7*5*2/1+3*9 = 100
 48.   4/8*6+7*5*2/1+9*3 = 100
 49.   6/2*1+5*7*8/4+3*9 = 100
 50.   6/2*1+5*7*8/4+9*3 = 100
 51.   6/2*1+5*8*7/4+3*9 = 100
 52.   6/2*1+5*8*7/4+9*3 = 100
 53.   6/2*1+7*5*8/4+3*9 = 100
 54.   6/2*1+7*5*8/4+9*3 = 100
 55.   6/2*1+7*8*5/4+3*9 = 100
 56.   6/2*1+7*8*5/4+9*3 = 100
 57.   6/2*1+8*5*7/4+3*9 = 100
 58.   6/2*1+8*5*7/4+9*3 = 100
 59.   6/2*1+8*7*5/4+3*9 = 100
 60.   6/2*1+8*7*5/4+9*3 = 100
 61.   6/2*9+5*7*8/4+1*3 = 100
 62.   6/2*9+5*7*8/4+3*1 = 100
 63.   6/2*9+5*8*7/4+1*3 = 100
 64.   6/2*9+5*8*7/4+3*1 = 100
 65.   6/2*9+7*5*8/4+1*3 = 100
 66.   6/2*9+7*5*8/4+3*1 = 100
 67.   6/2*9+7*8*5/4+1*3 = 100
 68.   6/2*9+7*8*5/4+3*1 = 100
 69.   6/2*9+8*5*7/4+1*3 = 100
 70.   6/2*9+8*5*7/4+3*1 = 100
 71.   6/2*9+8*7*5/4+1*3 = 100
 72.   6/2*9+8*7*5/4+3*1 = 100
 73.   6/8*4+2*5*7/1+3*9 = 100
 74.   6/8*4+2*5*7/1+9*3 = 100
 75.   6/8*4+2*7*5/1+3*9 = 100
 76.   6/8*4+2*7*5/1+9*3 = 100
 77.   6/8*4+5*2*7/1+3*9 = 100
 78.   6/8*4+5*2*7/1+9*3 = 100
 79.   6/8*4+5*7*2/1+3*9 = 100
 80.   6/8*4+5*7*2/1+9*3 = 100
 81.   6/8*4+7*2*5/1+3*9 = 100
 82.   6/8*4+7*2*5/1+9*3 = 100
 83.   6/8*4+7*5*2/1+3*9 = 100
 84.   6/8*4+7*5*2/1+9*3 = 100
 85.   7/1*9+2*3*5/6+4*8 = 100
 86.   7/1*9+2*3*5/6+8*4 = 100
 87.   7/1*9+2*5*3/6+4*8 = 100
 88.   7/1*9+2*5*3/6+8*4 = 100
 89.   7/1*9+3*2*5/6+4*8 = 100
 90.   7/1*9+3*2*5/6+8*4 = 100
 91.   7/1*9+3*5*2/6+4*8 = 100
 92.   7/1*9+3*5*2/6+8*4 = 100
 93.   7/1*9+5*2*3/6+4*8 = 100
 94.   7/1*9+5*2*3/6+8*4 = 100
 95.   7/1*9+5*3*2/6+4*8 = 100
 96.   7/1*9+5*3*2/6+8*4 = 100
 97.   7/4*8+3*6*9/2+1*5 = 100
 98.   7/4*8+3*6*9/2+5*1 = 100
 99.   7/4*8+3*9*6/2+1*5 = 100
100.   7/4*8+3*9*6/2+5*1 = 100
101.   7/4*8+6*3*9/2+1*5 = 100
102.   7/4*8+6*3*9/2+5*1 = 100
103.   7/4*8+6*9*3/2+1*5 = 100
104.   7/4*8+6*9*3/2+5*1 = 100
105.   7/4*8+9*3*6/2+1*5 = 100
106.   7/4*8+9*3*6/2+5*1 = 100
107.   7/4*8+9*6*3/2+1*5 = 100
108.   7/4*8+9*6*3/2+5*1 = 100
109.   8/1*4+2*3*5/6+7*9 = 100
110.   8/1*4+2*3*5/6+9*7 = 100
111.   8/1*4+2*5*3/6+7*9 = 100
112.   8/1*4+2*5*3/6+9*7 = 100
113.   8/1*4+3*2*5/6+7*9 = 100
114.   8/1*4+3*2*5/6+9*7 = 100
115.   8/1*4+3*5*2/6+7*9 = 100
116.   8/1*4+3*5*2/6+9*7 = 100
117.   8/1*4+5*2*3/6+7*9 = 100
118.   8/1*4+5*2*3/6+9*7 = 100
119.   8/1*4+5*3*2/6+7*9 = 100
120.   8/1*4+5*3*2/6+9*7 = 100
121.   8/3*9+4*5*7/2+1*6 = 100
122.   8/3*9+4*5*7/2+6*1 = 100
123.   8/3*9+4*7*5/2+1*6 = 100
124.   8/3*9+4*7*5/2+6*1 = 100
125.   8/3*9+5*4*7/2+1*6 = 100
126.   8/3*9+5*4*7/2+6*1 = 100
127.   8/3*9+5*7*4/2+1*6 = 100
128.   8/3*9+5*7*4/2+6*1 = 100
129.   8/3*9+7*4*5/2+1*6 = 100
130.   8/3*9+7*4*5/2+6*1 = 100
131.   8/3*9+7*5*4/2+1*6 = 100
132.   8/3*9+7*5*4/2+6*1 = 100
133.   8/4*7+3*6*9/2+1*5 = 100
134.   8/4*7+3*6*9/2+5*1 = 100
135.   8/4*7+3*9*6/2+1*5 = 100
136.   8/4*7+3*9*6/2+5*1 = 100
137.   8/4*7+6*3*9/2+1*5 = 100
138.   8/4*7+6*3*9/2+5*1 = 100
139.   8/4*7+6*9*3/2+1*5 = 100
140.   8/4*7+6*9*3/2+5*1 = 100
141.   8/4*7+9*3*6/2+1*5 = 100
142.   8/4*7+9*3*6/2+5*1 = 100
143.   8/4*7+9*6*3/2+1*5 = 100
144.   8/4*7+9*6*3/2+5*1 = 100
145.   9/1*7+2*3*5/6+4*8 = 100
146.   9/1*7+2*3*5/6+8*4 = 100
147.   9/1*7+2*5*3/6+4*8 = 100
148.   9/1*7+2*5*3/6+8*4 = 100
149.   9/1*7+3*2*5/6+4*8 = 100
150.   9/1*7+3*2*5/6+8*4 = 100
151.   9/1*7+3*5*2/6+4*8 = 100
152.   9/1*7+3*5*2/6+8*4 = 100
153.   9/1*7+5*2*3/6+4*8 = 100
154.   9/1*7+5*2*3/6+8*4 = 100
155.   9/1*7+5*3*2/6+4*8 = 100
156.   9/1*7+5*3*2/6+8*4 = 100
157.   9/2*6+5*7*8/4+1*3 = 100
158.   9/2*6+5*7*8/4+3*1 = 100
159.   9/2*6+5*8*7/4+1*3 = 100
160.   9/2*6+5*8*7/4+3*1 = 100
161.   9/2*6+7*5*8/4+1*3 = 100
162.   9/2*6+7*5*8/4+3*1 = 100
163.   9/2*6+7*8*5/4+1*3 = 100
164.   9/2*6+7*8*5/4+3*1 = 100
165.   9/2*6+8*5*7/4+1*3 = 100
166.   9/2*6+8*5*7/4+3*1 = 100
167.   9/2*6+8*7*5/4+1*3 = 100
168.   9/2*6+8*7*5/4+3*1 = 100
169.   9/3*8+4*5*7/2+1*6 = 100
170.   9/3*8+4*5*7/2+6*1 = 100
171.   9/3*8+4*7*5/2+1*6 = 100
172.   9/3*8+4*7*5/2+6*1 = 100
173.   9/3*8+5*4*7/2+1*6 = 100
174.   9/3*8+5*4*7/2+6*1 = 100
175.   9/3*8+5*7*4/2+1*6 = 100
176.   9/3*8+5*7*4/2+6*1 = 100
177.   9/3*8+7*4*5/2+1*6 = 100
178.   9/3*8+7*4*5/2+6*1 = 100
179.   9/3*8+7*5*4/2+1*6 = 100
180.   9/3*8+7*5*4/2+6*1 = 100
181.   9/6*4+2*5*7/1+3*8 = 100
182.   9/6*4+2*5*7/1+8*3 = 100
183.   9/6*4+2*7*5/1+3*8 = 100
184.   9/6*4+2*7*5/1+8*3 = 100
185.   9/6*4+5*2*7/1+3*8 = 100
186.   9/6*4+5*2*7/1+8*3 = 100
187.   9/6*4+5*7*2/1+3*8 = 100
188.   9/6*4+5*7*2/1+8*3 = 100
189.   9/6*4+7*2*5/1+3*8 = 100
190.   9/6*4+7*2*5/1+8*3 = 100
191.   9/6*4+7*5*2/1+3*8 = 100
192.   9/6*4+7*5*2/1+8*3 = 100

Edwin